Equivalent Units of Production and Related Costs

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process-Assembly Department
Bal., 5,000 units, 70% completed 19,500 To Finished Goods, 115,000 units ?
Direct materials, 118,000 units @ $1.80 212,400
Direct labor 265,200
Factory overhead 103,120
Bal. ? units, 45% completed ?

Determine the following:

a. The number of units in work in process inventory at the end of the period.
units

Feedback

Units in ending work in process represent units that have been started in the department but have not been completed, and therefore have not been transferred out to finished goods.

b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0".

Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole Units Equivalent Units Direct Materials Equivalent Units Conversion
Inventory in process, beginning
Started and completed
Transferred to finished goods
Inventory in process, ending
Total units

Feedback

When are materials added in the process? How does that affect the number of equivalent units in beginning work in process, units that have been started and completed and units in ending work in process? For units to be completed they have to be complete with respect to both materials and conversion.

c. Costs per equivalent unit for direct materials and conversion. If required, round your answers to the nearest cent.

Costs Per Equivalent Unit
Direct Materials $
Conversion $

d. Cost of the units started and completed during the period.
$

Analyzing changes in unit costs

The costs of materials consumed in producing good units in the Forming Department of Thomas Company were $5,250 and $14,670 for September and October, respectively. The number of equivalent units produced in September and October was 750 tons, and 1,630 tons, respectively. Evaluate the change in the cost of materials between the two months.

Round all answers to the nearest whole cent.

Energy cost per ton, September $
Energy cost per ton, October $

The cost of energy has appeared to increase  by $ per ton between September and October.
Answer #1

a. The number of units in work in process inventory at the end of the period.
5000+118000-115000 = 8000 units

b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0".

Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole Units Equivalent Units Direct Materials Equivalent Units Conversion
Inventory in process, beginning 5000 0 5000*30% = 1500
Started and completed 110000 110000 110000
Transferred to finished goods 115000 110000 111500
Inventory in process, ending 8000 8000

8000*45% = 3600

Total units 123000 118000 115100

c. Costs per equivalent unit for direct materials and conversion. If required, round your answers to the nearest cent.

Costs Per Equivalent Unit
Direct Materials $212400/118000 = 1.8
Conversion $368320/115100 = 3.2

d. Cost of the units started and completed during the period.
($1.8+3.2)*110000 = 550000

Evaluate the change in the cost of materials between the two months.

Round all answers to the nearest whole cent.

Energy cost per ton, September $5250/750 = 7
Energy cost per ton, October $14670/1630 = 9

The cost of energy has appeared to increase  by $2 per ton between September and October.